Home‎ > ‎Crew 133‎ > ‎

111-GreenHabReport

Greenhab Report,
Crew 133 
01/11/14

Prepared by XO Gartrelle
The Greenhab is operating normally.

The fan is operating normally set at 68 degrees F.

The heater was set to 40 degrees F at 9AM
The heater was set to 75 degrees F at 5PM

Only the hydroponic seedlings received water today.The onion, basil and cilantro seedlings have not germinated. The Mars Shoot is fine. The hydroponic seedlings were placed in the green hab at 9AM and returned to the upper Main hab for the night at 5PM. The lettuce seedlings have shown noticeable growth. The pepper seedlings have not sprouted but some are close to the soil surface. 

Tomorrow we will do a dry run of the hydroponic start up procedures. The purpose of the run through is to ensure we fully understand the procedure left by Greenhab Officer Ohrenstein from mission 132 and can execute it. Due to our water rationing, we will not use water and we will not actually turn on switches or open valves. This will be a complete simulation. Elizabeth Howell will assist me in this process. 

Photos from taken in the Greenhab today are attached. 

Greenhab readings for today are as follows:

                               Current              Max               Min   
Temperature         68 Degrees          91 Degrees       41 Degrees
Humidity              17%                    19%                 12%

Regards,
End Transmission